### v1

@@ VII Conclusions
This study provided insights into the influence of lower-limb joints on neural balance control in quiet stance and kneeling gaits in construction. The use of triple- and double-link inverted pendulum models, coupled with the intersection point height frequency analysis and the linear quadratic controller, allowed for a comprehensive evaluation of the balance strategies of the subject at the joint level. The integration of VR/MR tools to simulate elevated environments and welding tasks enabled the assessment of the impact of elevation and wearable knee exoskeletons on postural balance and task performance. The multi-subject experiment results highlight the critical role of knee joint strategies in maintaining postural stability and the potential of wearable exoskeletons in enhancing worker safety and performance in hazardous construction environments. The evaluation pipeline developed in this study can be further applied to assess the potential of wearable assistive devices across diverse occupational work settings.

### v2

@@ VII Conclusions
This study provided insights into the influence of lower-limb joints on neural balance control in quiet stance and kneeling gaits in construction. The use of triple- and double-link inverted pendulum models, coupled with the intersection point height frequency analysis and the linear quadratic controller, allowed for a comprehensive evaluation of the balance strategies of the subject at the joint level. The integration of VR/MR tools to simulate elevated environments and welding tasks enabled the assessment of the impact of elevation and wearable knee exoskeletons on postural balance and skilled task performance. The multi-subject experiment results highlight the critical role of knee joint strategies in maintaining postural stability and the potential of wearable exoskeletons in enhancing worker safety and performance in hazardous construction environments. The evaluation pipeline developed in this study can be further applied to assessing the potential of wearable assistive devices across diverse occupational work settings.